OpenStudy (anonymous):

the first term of a geometric series is -1, and the common ratio is -3. how many terms are in the series if its sum is 182?

OpenStudy (anonymous):

hi :)

OpenStudy (amistre64):

do you recall the formula for the sum of a geo series?

OpenStudy (amistre64):

\[S=a_1\frac{1-r^n}{1-r}\] solve for n \[S\frac{1-r}{a_1}=1-r^n\] \[S\frac{1-r}{a_1}-1=-r^n\] \[S\frac{r-1}{a_1}+1=r^n\]which becomes a log i bleive

OpenStudy (amistre64):

\[ln(S\frac{r-1}{a_1}+1)=ln(r^n)\] \[ln(S\frac{r-1}{a_1}+1)=n ln(r)\] \[\frac{ln(S\frac{r-1}{a_1}+1)}{ln~r}=n\]

OpenStudy (anonymous):

thnx

OpenStudy (amistre64):

yep, now its just a matter of filling in the values you are given and voila!! :)